Fermi National Laboratory
 

Accelerator Update for October 27, 2006 - October 30, 2006

 

Friday October 27

The day shift began with Operations monitoring the Tevatron (TeV) store 5030 (53.35E30) and the Recycler stash (122.5E10), with the Antiproton Source (Pbar) stacking (12.1E10), with Linac, Booster, and the Main Injector (MI) operating normally, with MiniBooNE and NuMI taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beams.

The Recycler completed stashing antiprotons at 11:40 AM.

Pbar experts began a study at 1:55 PM.

The Recycler completed stashing at 6:27 PM.

The Recycler completed stashing at 9:29 PM.

Operations terminated store 5030 at 10:02 PM.

Operations put the TeV into a shot setup at 10:15 PM.

Saturday October 28

The midnight shift began with the TeV in shot setup, with Operations monitoring the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam.

Operations established store 5031 at 12:35 AM with an initial luminosity of 176.97E30.

The Recycler completed stashing at 9:40 PM. This was the eighth and last transfer for the Saturday.

At 10:17 PM, the NuMI horn tripped off on a ground fault indication. It reset okay.

Sunday October 29

The midnight shift began with Operations monitoring store 5031 and the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam.

At 4:26 AM, a MI RF station (MIRF13) tripped off and wouldn’t reset.

Operations terminated store 5031 at 5:34 AM.

At 6:32 AM, the NuMI beam permit tripped off. Operators discovered that a +15 volt power supply for the BLM chassis had failed. Operations tried to contact experts.

Operations reported at 8 AM that the stack was at 92.8E10 and the stash was at 387.3E10.

At 8:21 AM, Operators reported that they were searching for spare BLM parts and for an expert to help with the repair. NuMI remained off due to this problem.

At 9:20 AM, Operations contacted an expert to help with the NuMI BLM chassis problem.

Operations put the TeV through a dry squeeze at 9:37 AM.

Operators found some spare BLM chassis cards. The expert arrived and with her help, Operations reestablished beam to NuMI at 10:48 AM.

Operations put the TeV into shot setup at 11:09 AM.

At 12:34 PM, the TeV quenched during the squeeze. Operators investigated.

TeV experts reported at 4:16 PM that the cause of the quench was related to the number of antiprotons. The TeV quenched at A1, BA, and BB. Operations loaded a 14x0 store (14 bunches of protons and no antiprotons) to study the problem, but no system failures occurred.

Operations put the TeV through a dry squeeze at 6:04 PM.

Operations put the TeV into shot setup at 9:57 PM.

Operations established store 5037 at 11:45 PM with an initial luminosity of 135.11E30.

The Recycler completed stashing at 11:57 PM; five transfers of antiprotons occurred on Sunday.

Monday October 30

The midnight shift began with Operations monitoring store 5037 and the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam.

The Recycler completed stashing at 5 AM.

At 7:34 AM, the TeV quenched at E4 due to an E4 cryo system failure. CHL technicians and operators investigated.

The Recycler completed stashing at 7:35 AM.

The Neutron Therapy Facility reported at 7:38 AM that they would be treating four patients this morning.

The Plans for Friday and the Weekend

The plan for today is to stash and stack. Later in the morning, techs will access the TeV F sector to investigate a possible vacuum valve problem. This access will shut off beam to Pbar, NuMI, and MiniBooNE. Pbar will transfer all its antiprotons before the F sector access and then access their tunnels to search for a LCW leak.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 10/30/2006    email Fermilab